OpenStudy (anonymous):

HELP ASAP MEDAL! Please show steps I would like to understand. Find the percent increase the following equation models? y=34(1.4)^x a. 1.4% b.34% c.40% d.140%

OpenStudy (anonymous):

@jdoe0001 do you mind helping?

OpenStudy (jdoe0001):

hmm do you know what they're asking for?

OpenStudy (anonymous):

how much it is increasing by?

OpenStudy (jdoe0001):

yes... I'm reading that too

OpenStudy (anonymous):

Well i don't know how to solve it haha

OpenStudy (jdoe0001):

hmmm not sure if we could get a fixed value for it though

OpenStudy (anonymous):

So what should I do?

OpenStudy (jdoe0001):

well ahemm so.... this is what I get from it, one sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\large y=34(1.4)^x\qquad \begin{array}{ccllll} x&y \\\hline\\ 1&34(1.4)^1\to 47.6\\ 2&34(1.4)^2\to 66.94\\ 3&34(1.4)^3\to 93.296\\ ...&... \end{array} \\ \quad \\ \textit{by how much did "y" increase from one spot to the next in }\%? \\ \quad \\ 66.94-47.6={\color{brown}{ 19.34}}\implies \begin{array}{llll} amount&\%\\ \\\hline\\ 47.6&100\\ {\color{brown}{ 19.34}}&x \end{array}\implies \cfrac{47.6}{{\color{brown}{ 19.34}}}=\cfrac{100}{x}\implies x=? \\ \quad \\ 93.296-66.94={\color{brown}{ 26.356}}\implies \begin{array}{llll} amount&\%\\ \\\hline\\ 66.94&100\\ {\color{brown}{ 26.356}}&x \end{array}\implies \cfrac{66.94}{{\color{brown}{ 26.356}}}=\cfrac{100}{x}\implies x=?\)

OpenStudy (anonymous):

Ugh...

OpenStudy (anonymous):

So i found out the answer is 40% but i cant find how it is/how they got that answer

OpenStudy (jdoe0001):

hehe, so there's a "difference" from one point to the next for "y" coordinate so seems to me you're asked what's that difference in approximate percent

OpenStudy (jdoe0001):

yeah 40% would be the approximate increase percentage

OpenStudy (anonymous):

but how do you know that?

OpenStudy (jdoe0001):

well..how did you get the 40% anyway?

OpenStudy (anonymous):

I didn't get it. I asked my friend but I can't figure out step by step how they got that

OpenStudy (jdoe0001):

well... check my posting, and solve those, maybe add more values to the table, say maybe x =4 and 5 and 6, and check \(y=34(1.4)^x\qquad \begin{array}{ccllll} x&y \\\hline\\ 1&34(1.4)^1\to 47.6\\ 2&34(1.4)^2\to 66.94\\ 3&34(1.4)^3\to 93.296\\ ...&... \end{array} \\ \quad \\ \textit{by how much did "y" increase from one spot to the next in }\%? \\ \quad \\ 66.94-47.6={\color{brown}{ 19.34}}\implies \begin{array}{llll} amount&\% \\\hline\\ 47.6&100\\ {\color{brown}{ 19.34}}&x \end{array}\implies \cfrac{47.6}{{\color{brown}{ 19.34}}}=\cfrac{100}{x}\implies x=? \\ \quad \\ 93.296-66.94={\color{brown}{ 26.356}}\implies \begin{array}{llll} amount&\% \\\hline\\ 66.94&100\\ {\color{brown}{ 26.356}}&x \end{array}\implies \cfrac{66.94}{{\color{brown}{ 26.356}}}=\cfrac{100}{x}\implies x=?\)

OpenStudy (jdoe0001):

seems in short, you just have to make a table of values, like above, for "x" and "y" and get the difference for the "y" values and then check by how much they went up, in percent

OpenStudy (jdoe0001):

for example say x y ----- 1 5 2 15 so... by how much "y" went up from 5 to 10? well it went up 15-5 = 10 so what's 10 in % of 5? well \(\begin{array}{llll} amount&\% \\\hline\\ 5&100\\ 10&x \end{array}\implies \cfrac{5}{10}=\cfrac{100}{x}\implies x=\cfrac{10\cdot 100}{5}\)